A person standing on a frictionless surface has a mass of 69.5kg including the gun is holding . He fires a 4.2g round at 955m/s.
Find the recoil velocity of the hunter if his fire horizontally.
1
Expert's answer
2021-09-28T16:54:19-0400

By the principle of conservation of impulse:

"m_1v_1+m_2v_2=0"

where m1 is mass of person, m2 is mass of bullet,

v1 is recoil velocity of the hunter, v2 is velocity of bullet.


"v_1=-\frac{m_2v_2}{m_1}=-\frac{4.2\cdot 10^{-3}\cdot 955}{69.5}=-0.06" m/s